OpenStudy (anonymous):

3x+4y=19 3x+6y=33 solve the system by elmination

OpenStudy (texaschic101):

3x + 4y = 19 3x + 6y = 33 -->(-1) ---------------- 3x + 4y = 19 -3x - 6y = -33 (result of multiplying by -1) -----------add -2y = - 14 y = 7 now sub 7 in for y in either equation to find x 3x + 4y = 19 3x + 4(7) = 19 3x + 28 = 19 3x = 19 - 28 3x = - 9 x = -3 check.. 3x + 6y = 33 3(-3) + 6(6) = 33 -9 + 42 = 33 33 = 33 (correct) x = -3 and y = 7

OpenStudy (anonymous):

Let's solve for y first. \[3x+4y=19\] \[3x+6y=33\] \[2y=14\] \[y=7\] Let's look at x \[3x+28=19\] \[3x=-9\] \[x=-3\] \[3x+42=33\] \[3x=-9\] \[x=-3\] Let's plug this back into the equations to check out work. \[3(-3)+4(7)=19 \] \[3(-3)+6(-7)=33\] It works! So, x=-3 and y=-7 I hope this helped!

OpenStudy (texaschic101):

oops..I messed up on the checking part 3x + 6y = 33 3(-3) + 6(7) = 33 -9 + 42 = 33 33 = 33 (correct) x = -3 and y = 7

OpenStudy (anonymous):

I meant 7, not -7. (Typo!!)

OpenStudy (texaschic101):

lol...a lot of typos here...but the answers are correct :)

OpenStudy (anonymous):

thankss